Solve by substitution method and what is the solution set?
9x+2y=-23
x=3-3y

Answer by mananth(16946)

9 x + 2 y = -23 .............1
3 y = 3 -1 x ---------2
/ 3 =
y = 1 -0.33 x
Substitute y in (1)
9 x + 2 ( 1 -0.33 x ) = -23
9 x + 2 + -0.67 x = -23
9 x -0.67 x = -23 -2
8.33 x = -25
/ 8.33
x= -3
Plug the value of x in (1)
9 x + 2 y = -23
9 * -3 + 2 y = -23
-27 + 2 y = -23
2 y = 4
/ 2
y= 2

CHECK
3 y = 3 + -1 x
6 = 3 + 3